Find the difference in entropy between one gram of ice 0°C and one gram of steam at 100°C. Latent heat of fission of ice is at 80 calories specific heat of water is 1 and latent heat of steam at 100°C with 540 calories
1
Expert's answer
2021-04-01T18:33:23-0400

Let the final temperature of the mixture will become T.

"Q_1=1\times 540+1\times 1\times(100-T)"

now energy required to convert ice to water

"Q_2=80+T"

Total energy always be conserve,

So,

"540+100-T=80+T"

"\Rightarrow 2T = 560"

"\Rightarrow T = 280^\circ C"

"\Delta S_1 = \frac{540}{273}"

"\Delta S_2 =-\frac{180}{273}"

"\Delta S_3=\frac{360}{273}"

"\Delta S=0"

Hence total change in entropy =0